A certain metal M forms a soluble sulfate salt M2(So4) Suppose the left half cell of a galvanic cell apparatus is filled with a 450. mM solution of M2(So4) and the right half cell with a 4.50 M solution of the same substance. Electrodes made of M are dipped into both solutions and a voltmeter is connected between them. The temperature of the apparatus is held constant at 30.0 cc. o left 10 Which electrode will be positive? right What voltage will the voltmeter show? Assume its positive lead is connected to the positive electrode Be sure your answer has a unit symbol, if necessary, and round it to 2 significant
Answer #1
which electride will be positive right
whate voltage will the volto meter shows 0.020V

Concentration of M3+ at anode = 0.450M × 2 = 0.900M

Concentratiom of M3+ at cathode = 4.50M × 2 = 9.00M

M(s) -------> M3+(aq) + 3e at anode

M3+(aq) (9.00M) + 3e ------> M(s) at cathode

number of electron transfer , n = 3

reaction quotient,Q= [M3+]low/[M3+]high

Q = 0.900M/9.00M = 0.1

Nernst equation for concentration cell at 25℃ is

Ecell = - (0.0592V/n)logQ

Ecell = - (0.0592V/3)log(0.1)

Ecell = 0.020V